OpenStudy (anonymous):

( 7- square root 54) - ( 5 + square root 24)

OpenStudy (jiteshmeghwal9):

\[7-\sqrt{54}-5-\sqrt{24}\] \[2-3\sqrt{2}-2\sqrt{6}\]

OpenStudy (jiteshmeghwal9):

Is it correct???

OpenStudy (anonymous):

can't it be simplified further?

OpenStudy (jiteshmeghwal9):

No.

OpenStudy (anonymous):

wouldn't it be 2 - square root 6? there aren't any options for the answer you came up with

OpenStudy (jiteshmeghwal9):

i have only solve the sqrt and subtract 7 with 2.

OpenStudy (jiteshmeghwal9):

Hey! wht were u typing???

OpenStudy (anonymous):

oh i was just about to say huh, I'm still a little confused o.O

OpenStudy (anonymous):

?????

OpenStudy (jiteshmeghwal9):

Wht is ur confusion???

OpenStudy (anonymous):

the answer you came up with isn't one of the options >,<

OpenStudy (anonymous):

@jiteshmeghwal9 \[\sqrt{54}=3\sqrt{2}\]????????

OpenStudy (anonymous):

its \[3\sqrt{6}\]

OpenStudy (anonymous):

look guys, the options are : 2- square root 78 2- square root 30 2- 5 square root 6 2- square root 6

OpenStudy (anonymous):

soo ??????????????????????????????????????????????????????????????

OpenStudy (anonymous):

????

OpenStudy (anonymous):

i need help!

OpenStudy (anonymous):

anyone PLEASE

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

\[2-5\sqrt{6}\]

OpenStudy (angela210793):